If \\(a=3i-2j+k\\), \\(b=2i-4j-3k\\) and \\(c=-i+2j+2k\\), find the magnitude of \\(a+b+c\\)
1
Expert's answer
2020-04-20T17:27:45-0400

Analytic Geometry

We need to find the magnitude of the sum of given vectors.

Solution:


The given vectors are,

"\vec a = 3i-2j+k\\
\vec b = 2i-4j-3k\\
\vec c =-i+2j+2k"

Sum of the vectors "= \vec a + \vec b + \vec c"

"= 3i-2j+k + 2i-4j-3k + -i+2j+2k"

"= 4i - 4 j"

Magnitude of the given vectors

"= |\vec a + \vec b + \vec c| = \sqrt {4^2 + (-4)^2} = \sqrt {(16 +16)} = \sqrt {32}"


"= 4 \sqrt 2"

Magnitude of the given vectors ="4 \sqrt 2"